# Locate-then-edit for Multi-hop Factual Recall under Knowledge Editing

### v2

@@ 4 IFMET: An Advanced Locate-then-Edit Method
Based on the previous key-value memories Geva et al. (2021), our method is based on the hypothesis that factual knowledge is stored within the Feedforward Neural Networks (FFNs) of MLPs. Specifically, for the $l$-th layer FFN of the $i$-th token, its output is given by: $v^{i}_{l}=f(W^{in}_{l}h^{i}_{l-1})W^{out}_{l}$, where $f(\cdot)$ is the activation function, and $h^{i}_{l-1}$ is the input of the $l$-th MLP layer (for simplicity, the superscript $l$ is omitted in the following discussion). In this context, $f(W^{in}h^{i})$ functions as the keys, denoted as $k_{i}$, the outputs of the subsequent layer represent the corresponding values $v_{i}$, and $W^{out}$ denotes the weights of the knowledge stored in the FFN that needs modifying. Such a structure is well aligned with the triplet form in a fact $(s,r,o)$, where the keys $k_{i}$ correspond to entities of interest $s_{i}$ or some specific fact $(s_{i},r_{i})$ and values $v_{i}$ contain information about $o_{i}$. Thus, we have $W^{out}k=v$ for $(k,v)$, which represents the fact $\left(s,r,o\right)$ (Geva et al., 2021). We aim to modify $W^{out}$ such that $W^{out}k=v^{*}$, where $v^{*}$ contains the information of the new knowledge.
@@
Motivated by the above, in IFMET, considering a modification, there are two steps for both earlier and latter layers in MLPs: $Search$ and $Calculate$. The $Search$ process identifies the suitable $v^{*}$ through the edit prompt corresponding to the triplet. Then the $Calculate$ process computes the change in weights $W^{out}$ using $v^{*}$. These two processes are foundational in existing knowledge editing methodologies. In experiments, we adopt the state-of-the-art locate-then-edit method PMET (Li et al., 2024c). The primary differences between the first and further edits are reflected in the edit prompt and the layers edited. Specifically, for the edit instance $e=(s,r,o\to o*)$, the first edit utilized a one-hop edit template $T_{r}(s)$ provided by MQuAKE to edit early layers of the model. For the furtherance edit, a two-hop template $T_{C}(s^{\prime})$ composed of a support case $(s^{\prime},r,s)$ and $(s,r,o^{*})$ was used, and this template was applied to edit later layers of the model. Due to space limitations, the flowchart of the algorithm and related implementation details are provided in Algorithm 1 and Appendix  C.
@@ 5.1 Experimental Setup
Dataset. We use MQuAKE-3K (Zhong et al., 2023), a challenging dataset designed to evaluate models’ ability to perform multi-hop reasoning with newly edited knowledge. Each entry consists of multiple single-hop edits and includes multi-hop reasoning questions, placing heightened demands on the ability of edited models to effectively utilize the updated information.
Baselines. As IFMET is a locate-then-edit approach, we mainly compare it with previous weight-modifying approaches. Specifically, our baseline includes the following methods: Original, which refers to the original model without any edits; ROME Meng et al. (2022a), which identifies editing areas using causal mediation analysis framed as a least-squares problem under linear equality constraints and solving it using Lagrange multipliers; MEND Mitchell et al. (2022), which employs meta-learning to train a hypernetwork for inferring weight updates from gradients; MEMIT Meng et al. (2023), which extends ROME to edit a large set of facts by updating weights in a range of layers; MeLLo, which manages multi-hop knowledge editing by decomposing subproblems and detecting conflicts; PMET, which optimizes FFN hidden states for precise weight updates, achieving SOTA performance in COUNTERFACT (Meng et al., 2022a) and ZsRE (Levy et al., 2017).
Setup and Hyperparameters: To evaluate the performance, we employ Multi-hop Accuracy (Zhong et al., 2023). Our experiments are conducted on the GPT-J (6B) model. We use PMET as our primary experimental method for both the first and furtherance edits. We construct a supplementary set from the knowledge triples of MQuAKE-3K to support our IFMET. When constructing the support set, for each edit case, no more than three supplements per relation were added from the supplementary dataset. The relation types of the supplementary set are the same as MQuAKE. We set the edit batch sizes to 1, 1000, and 3000. Additional details are presented in Appendix D.2.
@@ 5.2 Experimental Results
Model Method Batch_size=1 Batch_size=1000 Batch_size=3000
GPT-J-6B Original 39.63 - -
MeLLo  Zhong et al. (2023) 20.3 11.0 10.2
ROME  Meng et al. (2022a) 7.6 - -
MEMIT Meng et al. (2023) 12.3 8.1 1.8
MEND  Mitchell et al. (2022) 11.5 4.3 3.5
PMET  Li et al. (2024c) 11.17 11.13 11.7
IFMET (ours) 23.04 18.8 16.07
Table 3: Performance comparison of different methods on the MQuAKE-CF dataset with few-shot. Methods with indicate weight-modifying methods, while methods with are weight-preserving methods. ‘-’ indicates no relevant result, as ROME does not support multiple edits.
General performance. Table 3 demonstrates the performance of various established methods alongside IFMET on MQuAKE-3K. We can easily see the previous weight-modifying approaches generally exhibited poor performance. As the edit batch size increases, all methods except PMET show a certain downward trend. Our method inherits the good batch editing ability of PMET and consistently outperforms all others, showcasing a leading edge. Our approach significantly improves upon existing knowledge editing techniques, demonstrating the effectiveness and necessity of updating knowledge storage in deeper MLP layers.
Editor Correct Answer $\uparrow$ Original Answer $\downarrow$
Average Accuracy Pre Mid Post Average Accuracy Pre Mid Post
GPT-J 7.70 6.03 16.92 7.00 39.63 38.43 35.9 44.27
GPT-J+CoT 6.83 5.92 9.23 7.76 42.83 41.56 39.74 47.33
PMET 11.17 12.13 16.09 6.52 29.95 23.60 35.66 41.85
PMET+CoT 17.04 19.84 14.32 11.91 29.35 23.12 30.43 43.22
IFMET 23.04 20.24 15.28 33.38 23.08 20.18 34.32 24.25
IFMET+CoT 31.01 31.69 19.49 35.15 21.62 17.71 30.51 26.27
Table 4: Performance comparison between correct and original answers using PMET and our editors, with an edit batch size of 1. The type of edited fact—Pre, Mid, or Post—depends on the edited data position within the multi-hop reasoning chain. Average accuracy is calculated as the weighted average of results from these three categories, which have respective quantities of 1824, 390, and 786. Additionally, +CoT denoted the performance incorporating a Chain-of-thought (CoT) prompt, which can significantly enhance model performance for reasoning tasks.
Effect of number of hops. Table 5 in the Appendix displays the performance trends of various knowledge editing methods with different numbers of hops in multi-hop factual recall. We can see that each additional reason hop will negatively impact performance. Notably, IFMET is the best one in all cases, with minimal performance degradation. In particular, its results are close to those of the original model in the two-hop scenarios. This slight decrease underlines IFMET’s robustness and its superior ability to handle complex multi-hop tasks effectively.
Effect of the number of edited instances. We then consider the performance with different edited instances, which refer to the required number of new knowledge updates in the edit case. The results shown in Appendix Table 6 indicate a performance decline across all methods as more edits are introduced. Notably, IFMET consistently outperforms other approaches, showing the smallest average decline across different instance scenarios. Surprisingly, IFMET achieves an accuracy distribution that is close to that of the original model. It is also the only method that maintains excellent performance in complex two-hop and four-hop scenarios, which even outperforms single-hop cases. This can be more visually observed in Appendix Figure 6.
Effect of the edit position. As previously mentioned, the same single-hop fact requires different layers to provide knowledge, depending on its position in the multi-hop reasoning chain, involving the earlier and later MLPs. By categorizing according to position, we can assess whether the editing methods have comprehensively updated the relevant knowledge in the model rather than just making partial updates. We classify the edited case according to its position in the relevant multi-hop reasoning chain as Pre, Mid, and Post. For instance, in a three-hop knowledge sequence, editing the first hop is classified as pre, the second as mid, and the third as post. Please refer to Appendix D.1 for the classification of more complex, multi-edit scenarios. To assess the completeness of our method, we evaluated its performance for both eliminating original knowledge and incorporating new knowledge. As detailed in Table 4, our method significantly enhances outcomes across all classification types—Pre, Mid, and Post. Notably, it achieves exceptional improvements in both modifying new knowledge and eliminating original knowledge, especially in cases classified as Post.
@@ Appendix A Related Work
Mechanistic Interpretability serves as a tool for debugging and enhancing LLMs and can be applied to a variety of downstream tasks. Xiao et al. (2024) leveraged explanations from multi-head self-attention (MHSA) mechanisms in LLMs by introducing StreamingLLM, a model capable of handling unlimited text without requiring fine-tuning. Through causal tracing, Hendel et al. (2023); Todd et al. (2024) demonstrated that certain attention heads can efficiently encode compact representations of example tasks, leading to improved performance in few-shot prompting. Liu et al. (2024) explored the role of social bias in LLMs, introducing the concept of social bias neurons to explain and mitigate such biases. Furthermore, Li et al. (2024b) proposed an intervention technique during inference, which, based on the interpretability of attention heads, shifts activation values toward “truthful” responses to reduce model hallucinations. In this paper, we analyze the MLP and MHSA components of LLMs to uncover the mechanisms that enable multi-hop reasoning,and building on our findings, we introduce a targeted knowledge-editing method IFMET.
Appendix B More Details
B.1 Subset of MQuAKE
B.1.1 1-hop and 2-hop subset for Mechanism Exploration
In exploring the mechanisms of fact recall for one-hop and two-hop queries, this experiment utilized cloze templates as the experimental framework. We extracted knowledge from MQuAKE that could answer cloze templates in a zero-shot setting. This approach ensured that the model could recall the knowledge under the strictest conditions while minimizing the impact of unclear responses on the experimental results. The distribution of various relation types across the two subsets is illustrated in Figure 4.
@@ B.1 Subset of MQuAKE
B.1.2 Pre and Post Subset
To construct the subset, we selected two-hop queries from MQuAKE with Cloze-Format templates, and then randomly drew a nearly equal number($\approx 300$) of cases based on the proportion of relations.
B.2 Least squares and Minimum-norm method
@@
To minimize the probability of the intermediate answer $j$, we replace its logits with the smallest logits of the model’s vocabulary, and provide appropriate compensation for the final answer $k$ to maintain the probability of the final answer unchanged. The $\Delta h$ can be represented as:
@@
The change in the probability of the final answer after causal intervention can be represented by the function $f(\alpha)$: $f(\alpha)=P(h^{*},k)-P(h,k)$ Where $f(\alpha)$ is a monotonically increasing function on the interval $(0,1)$. We can find the zero of this function using the bisection method, ensuring that the final answer, after causal intervention, remains within an acceptable error margin with unchanged probability.
B.3 Causal Intervention on Single-hop Case
@@
The results of intervention for single-hop cases are shown in Figure 5. Except for the input layer, no significant effects are shown, indicating that in the single-hop fact recall task, the prediction of the final answer at the last token position is largely independent of the information from the intermediate results.
Appendix C Details of IFMET
C.1 Detailed Supplementary Set Construction Process
We collect 2615 subjects from the MQuAKE dataset. For each subject $s$, we use a Wikidata SPARQL query to retrieve the triplet $(s^{\prime},r^{\prime},s)$. The query is illustrated in Table 7. To keep the query complexity within an acceptable range, we collected all relationships that have appeared in MQuAKE and restricted $r^{\prime}$ to those that have occurred in the relation set. We then use the prompt 9 to filter out the answerable $(s^{\prime},r^{\prime},s)$ triples. For each edit case $(s,r,o\to o*)$, we are able to construct a two-hop edit template $T_{C}(s^{\prime})$ with the multi-hop chain $C=(s^{\prime},r^{\prime}s)\oplus(s,r,o\to o*)$.
C.2 Detailed Edit Process

### v3

@@ 4 IFMET
Based on previous key-value memories (Geva et al., 2021), our method to edit the MLP is based on the hypothesis that factual knowledge is stored within the Feedforward Neural Networks (FFNs) of MLPs. Specifically, for the $l$-th layer FFN, its output of the $i$-th token’s hidden state $h^{i}_{l-1}$, is given by: $v^{i}_{l}=f(W^{in}_{l}h^{i}_{l-1})W^{out}_{l}$, where $f(\cdot)$ is the activation function, and $h^{i}_{l-1}$ is the input of the $l$-th MLP layer (for simplicity, the superscript $l$ is omitted in the following discussion). In this context, $f(W^{in}h^{i})$ functions as the keys, denoted as $k_{i}$, the outputs represent the corresponding values $v_{i}$, and $W^{out}$ denotes the weights of the knowledge stored in the FFN that needs modifying. Such a structure is well aligned with the triplet form $(s,r,o)$, where the keys $k_{i}$ correspond to entities of interest $s_{i}$ or some specific fact $(s_{i},r_{i})$ and values $v_{i}$ contain information about $o_{i}$. Thus, we have $W^{out}k=v$ for $(k,v)$, which represents the fact $\left(s,r,o\right)$ (Geva et al., 2021). We aim to modify $W^{out}$ such that $W^{out}k=v^{*}$, where $v^{*}$ contains the information of the new knowledge.
@@
Motivated by the above, in IFMET, there are two main steps for both the first and second edit stages: $Search$ and $Calculate$. The $Search$ process identifies the suitable $v^{*}$ through the edit prompt. Then the $Calculate$ process computes the change in weights $W^{out}$ using $v^{*}$. These two processes are foundational in existing knowledge editing methodologies. In experiments, we adopt the state-of-the-art locate-then-edit method PMET (Li et al., 2024c). The primary differences between the first and further edit stages are reflected in the used edit prompt and the layers edited. Specifically, for the edit instance $e=(s,r,o\to o*)$, the first edit utilized a single-hop edit prompt $T_{e}$ provided by the dataset to edit shallow layers of the model. For the furtherance edit, the two-hop prompt $T_{C_{post}}$ composed of $(s^{\prime},r,s)$ and $(s,r,o^{*})$ was used, and this prompt was applied to edit deeper layers of the model. Due to space limitations, the flowchart of the algorithm and related implementation details are provided in Algorithm 1 and Appendix D.2.
@@ 5.1 Experimental Setup
Dataset and Baselines [^4] More details about the dataset in Appendix B.1. We mainly compare IFMET with previous weight-modifying approaches, especially these single-stage edit methods applying shallow MLP edits based on single-hop edit prompts. Comparison with weight-preserving methods is discussed in section G.1. MQuAKE-3K (Zhong et al., 2023), a challenging and widely used dataset designed to evaluate models’ ability to perform multi-hop fact recall with newly edited knowledge. Each edit instance contains a multi-hop factual recall chain and the corresponding multi-hop textual question. The instance includes at least one fact from the chain to be edited and provides its single-hop edit prompt. Baselines are Base, which refers to the original GPT-J(6B) model without any edits; FT basic fine-tuning method; MEND (Mitchell et al., 2022), which employs meta-learning for weight updating; ROME (Meng et al., 2022a), the classic single-stage locate-then-edit method; MEMIT (Meng et al., 2023), which extends ROME to edit a large set of facts by updating weights in a range of layers; PMET, locate-then-edit method with FFN optimization.
Setup and Hyperparameters. To evaluate the performance of different KE methods, we adopt Multi-hop question answering accuracy(Multi-hop Acc) as the primary metric. For each query, the unedited answer denotes the expected old fact before knowledge editing, while the edited answer represents the expected new fact after editing. We use PMET as our primary experimental method for both the first and furtherance edits and construct multi-hop edit prompts from the knowledge triples of MQuAKE-3K to support our IFMET. Additional details are presented in Appendix E.2.
[…]
Base 7.70 6.03 16.92 7.00
Base+CoT 6.83 5.92 9.23 7.76
PMET 11.17 12.13 16.09 6.52
PMET+CoT 17.04 19.84 14.32 11.91
IFMET 23.04 20.24 15.28 33.38
IFMET+CoT 31.01 31.69 19.49 35.15
(a) Edited Answer
Editor Average $\downarrow$ Pre $\downarrow$ Mid $\downarrow$ Post $\downarrow$
Base 39.63 38.43 35.9 44.27
Base+CoT 42.83 41.56 39.74 47.33
PMET 29.95 23.60 35.66 41.85
PMET+CoT 29.35 23.12 30.43 43.22
IFMET 23.08 20.18 34.32 24.25
IFMET+CoT 21.32 17.71 28.20 26.27
(b) Unedited Answer
Table 4: Multi-hop accuracy comparison for edited and unedited answers using PMET and our editors on the MQuAKE-3K dataset. Average accuracy is calculated as the weighted average of results from these three categories, which have respective quantities of 1824, 390, and 786. Additionally, +CoT denoted the performance incorporating a Chain-of-thought (CoT) prompt.
Editor Multi-hop $\uparrow$ Efficacy $\uparrow$
IFMET 28.38 ($\uparrow$78.0%) 99.56 ($\uparrow$12.8%)
w/o $First$ 23.14 ($\uparrow$45.1%) 66.59 ($\downarrow$24.6%)
w/o $Multi$ 17.69 ($\uparrow$10.9%) 100.00 ($\uparrow$13.3%)
w/o $Deeper$ 15.07 ($\downarrow$5.4%) 99.56 ($\uparrow$12.8%)
PMET 15.94 88.21
Table 5: The results of the ablation experiments on GPT-J-6B model using a subset of MQuAKE-CF. Both the percentages of decrease($\downarrow$) and increase($\uparrow$) are calculated relative to PMET as the baseline. The most significant performance decline is highlighted in red and the most significant performance increase is highlighted in green.
@@ 5.2 Experimental Results
General performance. Table 3 demonstrates the performance of various established methods alongside IFMET on MQuAKE-3K. To thoroughly explore the model’s ability to leverage the newly edited knowledge, we use Chain-of-Thought (CoT) prompting to guide the model’s responses to multi-hop tasks in this experiment. # Hops refers to the number of hops of the multi-hop fact chain in the edit instance, with a maximum of 4 hops. # Edits quantifies how many individual facts within the chain are edited, and its maximum value is the same as the maximum number of hops in the instance. What can be observed is that, on the overall average performance, IFMET consistently outperforms previous methods by a significant margin. Moreover, the performance improvement of IFMET is consistent across all subsets (e.g., 2-edit or 2-hop) of the dataset. Notably, in more complex reasoning scenarios, such as when edits>2 or hops>3, IFMET achieves a performance improvement of two to three times. This demonstrates the adequacy of knowledge updates in the IFMET method.
Comparison between the unedited answer and edited answer. In Section 3.2, we attempted to explain and demonstrate why existing methods still tend to output the unedited answer in multi-hop tasks, especially in the post-type multi-hop fact recall task. Here, we provide a more detailed breakdown of the dataset to investigate whether IFMET effectively alleviates this issue. Please refer to Appendix E.1 for the classification of {Pre,Mid,Post}. Results in Table 4 show that, in the Mid and Post scenarios where the facts are typically treated as implicit reasoning steps, IFMET effectively reduces the output of the unedited answer and improves the accuracy of the correct answer. Even in the pre-type tasks, where previous methods are relatively more proficient, IFMET achieves a significant improvement.
Ablation study. With Efficacy metric, which measures whether the model can successfully answer the single-hop fact recall prompt, we comprehensively evaluate the model’s ability to perform both single-hop and multi-hop reasoning using the edited knowledge. The importance of each component is reflected through comparisons of performance improvements over PMET. From the analysis of the ablation experiments in Table 5, we derive the following conclusions: w/o $First$: Only modifying the deeper layers using second stage edit with multi-hop edit prompt effectively enhances performance on multi-hop reasoning tasks. However, the absence of the single-hop edit prompt in the first stage resulted in the shallow MLP layers not being updated, leading to poor performance in single-hop fact recall tasks. This highlights the importance of the two-stage editing process. w/o $Multi$: In the second editing stage, we try to use the original single-hop edit prompt instead of a multi-hop edit prompt to edit the deeper MLP layers. However, the results corroborate our interpretability analysis which emphasizes the differences between single-hop and multi-hop reasoning mechanisms. Single-hop prompts cannot correctly modify knowledge in deep MLPs, highlighting the critical importance of multi-hop prompts. w/o $Deeper$: In this setup, we try to use the multi-hop edit prompt to edit shallow MLP layers (rather than deeper MLP layers). As observed across the table, there was a consistent minor fluctuation in performance. In contrast to IFMET’s +70% improvement, this underscores the necessity of editing knowledge in the deeper MLP layers when using the multi-hop prompts. More comprehensive ablation studies and discussions can be found in Appendix F.
Generalizability. Additionally, we conducted experiments on larger edit batches, newer models and more metrics in Appendix G, all of which showed significant performance improvements, effectively demonstrating the generalizability of our approach.
@@ 6 Conclusion
Impact Statement
This paper presents work whose goal is to advance the field of knowledge editing. We aim to enhance the locate-then-edit paradigm to address multi-hop factual recall tasks. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ Appendix A Related Work
Mechanistic Interpretability serves as a tool for debugging and enhancing LLMs and can be applied to a variety of downstream tasks. Xiao et al. (2024) leveraged explanations from multi-head self-attention (MHSA) mechanisms in LLMs by introducing StreamingLLM, a model capable of handling unlimited text without requiring fine-tuning. Through causal tracing, Hendel et al. (2023); Todd et al. (2024) demonstrated that certain attention heads can efficiently encode compact representations of example tasks, leading to improved performance in few-shot prompting. Liu et al. (2024) explored the role of social bias in LLMs, introducing the concept of social bias neurons to explain and mitigate such biases. Furthermore, Li et al. (2024b) proposed an intervention technique during inference, which, based on the interpretability of attention heads, shifts activation values toward “truthful” responses to reduce model hallucinations. In this paper, we analyze the MLP and MHSA components of LLMs to uncover the mechanisms that enable multi-hop reasoning and, building on our findings, we introduce a targeted knowledge-editing method, IFMET.
Appendix B Dataset
B.1 Details of MQuAKE-CF-3K
[…]
B.2 Subset of MQuAKE
B.2.1 1-hop and 2-hop subset for Mechanism Exploration
In exploring the mechanisms of fact recall for one-hop and two-hop queries, this experiment utilized cloze templates as the experimental prompts. We extracted knowledge from MQuAKE that could be answered by GPT-J-6B in a zero-shot setting. This approach ensured that the model could recall the knowledge under the strictest conditions while minimizing the impact of unclear responses on the experimental results. The distribution of various relation types across the two subsets is illustrated in Figure 4.
@@ B.2 Subset of MQuAKE
B.2.2 Pre and Post Subset
To construct the subset, we selected two-hop queries from MQuAKE with Cloze-Format templates and then randomly drew a nearly equal number($\approx 300$) of cases based on the proportion of relations.
Appendix C Causal Intervention
C.1 Least squares and Minimum-norm method
@@
To minimize the probability of the intermediate answer $j$, we replace its logits with the smallest logits of the model’s vocabulary and provide appropriate compensation for the final answer $k$ to maintain the probability of the final answer unchanged. The $\Delta h$ can be represented as:
@@
The change in the probability of the final answer after causal intervention can be represented by the function $f(\alpha)$: $f(\alpha)=P(h^{*},k)-P(h,k)$ Where $f(\alpha)$ is a monotonically increasing function on the interval $(0,1)$. We can find the zero of this function using the bisection method, ensuring that the final answer, after the causal intervention, remains within an acceptable error margin with unchanged probability.
C.2 Causal Intervention on Single-hop Case
@@
The results of intervention for single-hop cases are shown in Figure 5. Except for the input layer, no significant effects are shown, indicating that in the single-hop fact recall task, the prediction of the final answer at the last token position is largely independent of the information from the intermediate subjects.
C.3 Comparative experiments on the same fact at different positions in multi-hop factual recall
We sampled two subsets with approximately equal sizes from the MQuAKE-3K dataset, detailed in Appendix  B.2.2. By applying the SOTA locate-then-edit method PMET to layer [3-8], which follows (Li et al., 2024c), we present the percentage of cases where both pre-edited and post-edited models answer successfully in QA format or Cloze format under different edit batches. Notably we show the performance of pre-edited model on unedited answer, as well as the edited answer for the posted-edited model.
Appendix D Details of IFMET
D.1 Detailed Multi-hop Edit Prompts Construction Process
We provide two methods for constructing multi-hop edit prompts. The first method is used in the main experiments of this paper, while the performance of multi-hop edit prompts generated by the second method is discussed in the appendix G.
WikiData as Knowledge Base Practically, we utilize WikiData555www.wikidata.org to construct. We start by extracting all 2615 subjects from the MQuAKE dataset’s edits and deduplicating them to form a set of subjects $S_{e}=\{s_{i}|i=1,\dots\}$. For each subject $s$, We then perform a WikiData SPARQL query [^6] https://query.wikidata.org/ to identify a set of triplets for each subject $s_{i}$: $Sup=\{(s^{\prime},r^{\prime},o^{\prime})|o^{\prime}=s_{i}\}$. The query is illustrated in Table 14. To keep the query complexity within an acceptable range, we collected all relationships that have appeared in MQuAKE and restricted $r^{\prime}$ to those that have occurred in the relation set. To ensure the reliability of these facts, we then use the prompt 11 to filter out the answerable $(s^{\prime},r^{\prime},s)$ triples. For each edit case $(s,r,o\to o*)$, we are able to construct a two-hop edit template $T_{C}(s^{\prime})$ with the multi-hop chain $C=(s^{\prime},r^{\prime}s)\oplus(s,r,o\to o*)$.
Model as Knowledge Base We used a simple prompt in table 13 to retrieve relevant knowledge directly from the model to be edited for constructing the multi-hop edit prompt. Due to computational and time constraints, we limited each case to a minimum of one multi-hop edit prompt and a maximum of five multi-hop edit prompts.
D.2 Detailed Edit Process